A person is standing 400m from the base of a building that has a flag pole on top. If the angle of elevation from the person to the top of the building is 27.3 degrees, and the angle of elevation from the person to the top of the flag pole is 29.8 degrees, what is the height of the flag pole?
Answer by edjones(7311)

tan=opposite/adjacent
tan(28.9)=x/400
.55203=x/400
x=400*.55203
=220.812... m
.
tan(27.3)=x/400
.516138=x/400
x=400*.516138
=206.455... m
.
220.812-206.455
=14.3565... m
.
Ed